Carl would like to save $100,000 by his 40th birthday to pay for a special midlife crisis vacation. He plans to achieve this by investing equal annual amounts each year beginning on his 24th birthday and ending and including a payment on his 40th birthday. If the investment pays an 11 percent interest rate, what is the size of each annual payment Carl needs to invest?
Answer #1

Contribution is made from 24 th birthday to 40th birthday.

It means number of annual payments = 17

and payment is from beginning of year

Interest rate = 11% or 0.11

Future value of annuity formula = P *{ (1+r)^n - 1 } / r

100000 = P *((1+0.11)^17 - 1)/0.11

100000 = P *44.50084281
P = 100000 /44.50084281
P =2247.14845   

So, size of each equal annual payment required is $2247.15.

Please thumbs up